Based on a survey of 130 AUS students, we found 36 of them are using iPhone. Do we have enough evidence to conclude that about 30% of AUS students are using iPhone at 15% level of significance?
1
Expert's answer
2020-12-28T18:34:36-0500

At 15% level of significance implies an 85% confidence level. We create a 85% Confidence interval

85% CI="\hat p±z^* \sqrt{\frac{\hat p(1- \hat p)}{n}}"

"\hat p=\frac {36}{130}=0.2769"

="0.2769±1.44\sqrt{\frac{0.2769(1-0.2769)}{130}}"

=(0. 2204, 0.3334)

Since 0.3 is within the range, we can conclude that about 30%of AUS students are using iPhone at 15 %level of significance.